Problem 9. (20 points) If X is a normal random variable with parameters μ and σ2-49, compute (a) Píx> 5H (b) P(4<X < 16) (c) P(X < 8 (d) P(x < 20); (e) P(x> 16). 12
Answer #1

Ans:

Given that

mean=12

and

standard deviation=sqrt(49)=7

a)

z=(5-12)/7

z=-1

P(z>-1)=0.8413

b)

z(4)=(4-12)/7=-1.14

z(16)=(16-12)/7=0.57

P(-1.14<z<0.57)=P(z<0.57)-P(z<-1.14)

=0.7156-0.1271=0.5885

c)

z=(8-12)/7=-0.57

P(z<-0.57)=0.2843

d)

z=(20-12)/7=1.14

P(z<1.14)=0.8729

e)

z=(16-12)/7=0.57

P(z>0.57)=0.2843